1/6|4x-8|=6x what's the first thing that I do?
Answer by rapaljer(4671)

I would multiply both sides of the equation by the denominator, which is 6, to clear the fraction.
|4x-8| = 36x

That is the FIRST thing I would do. Did you by any chance want to see the second and third thing to do, or was that your only question?

Next, remember that there are two solutions for an absolute value problem like this. Also, you will need to check the answers to make sure you do not accidentally get an absolute value equal to a negative. If this happens you must reject the answer. You must solve:

FIRST SOLUTION:
4x-8 = 36x

Subtract 4x from each side:
-8 = 32x

Divide both sides by 32:
x= -8/32 = -1/4

This answer must be rejected, since in the original problem, it causes an abolute value to equal a negative, which is not allowed.

SECOND SOLUTION:
4x-8 = -36x

Subtract 4x from each side:
-8=-40x

Divide both sides by -40
x=-8/-40 = 1/5

The final answer should be x= 1/5 or x = 0.2

R^2 at SCC